# High Dimensional Distributed Gradient Descent with Arbitrary Number of Byzantine Attackers

### v2

@@ Appendix C Proof of Theorem 3
$\text{P}(\hat{Z}\neq Z)$ $\overset{(a)}{\geq}$ $\frac{H(Z|\mathbf{Y}^{S_{0}},\mathbf{X}^{A})-\ln 2}{\ln(d-1)}$ (113)
$=$ $\frac{H(Z)-I(Z;\mathbf{Y}^{S_{0}},\mathbf{X}^{A})-\ln 2}{\ln(d-1)}$
$\geq$ $1-\frac{I(Z;\mathbf{Y}^{S_{0}},\mathbf{X}^{A})+\ln 2}{\ln d}$
=$\overset{(b)}{=}$ $1-\frac{I(Z;\mathbf{X}^{A})+\ln 2}{\ln d}.$
In (a), $\mathbf{Y}^{S_{0}}$ means $\mathbf{Y}_{i}$ for all $i\in S_{0}$. For (b), note that the distribution of $\mathbf{Y}$ does not depend on $Z$, hence $I(Z;\mathbf{Y}^{S_{0}},\mathbf{X}^{A})=I(Z;\mathbf{X}^{A})$. The mutual information can be bounded by
$I(Z;\mathbf{X}^{A})=N_{A}I(Z;\mathbf{X})=N_{A}(H(\mathbf{X})-H(\mathbf{X}|Z)).$ (114)
For any $i$, given $Z=i$, $X_{i}$ is fixed. Therefore
$H(\mathbf{X})-H(\mathbf{X}|Z=i)$$\leq$ $-2\beta\ln\beta-(1-2\beta)\ln(1-2\beta)$ (115)
@@
$\text{P}(\hat{Z}\neq Z)\geq 1-\frac{2N_{A}\beta\left(\ln\frac{1}{\beta}+1\right)+\ln 2}{\ln d}.$ (116)
Now we restrict $N_{A}$ to ensure that $\text{P}(\hat{Z}\neq Z)\geq 1/2$. This yields
$N_{A}\leq\frac{\ln\frac{d}{4}}{4\beta\left(\ln\frac{1}{\beta}+1\right)}.$ (117)
Recall (111). For $i,j\in[d]$, $i\neq j$, $\|\mu_{i}-\mu_{j}\|_{2}\geq\sigma/\sqrt{\beta}$. Therefore, if $\hat{Z}\neq Z$, then
$\|\hat{\mu}-\mu^{*}\|_{2}$ =$\overset{(a)}{=}$ $\|\hat{\mu}-\mu_{Z}\|_{2}$ (118)
@@
$\geq$ $\frac{\sigma}{2\sqrt{\beta}}.$
For (a), note that given $Z=i$, the real mean of $\mathbf{X}$ is $\mu^{*}=mu_{i}$. For (b) uses (112), which yields $\|\hat{\mu}-\mu_{\hat{Z}}\|_{2}\leq\|\hat{\mu}-\mu_{Z}\|_{2}$.
From the above analysis, we conclude that if (117) holds, then with probability at least $1/2$, $\|\hat{\mu}-\mu^{*}\|_{2}\geq\sigma/(2\sqrt{\beta})$. Taking $\beta=\alpha/2$. From Lemma 11, with probability at least $1/2-\exp[-(\ln 2-1/2)N\alpha]$,
$\|\hat{\mu}-\mu^{*}\|_{2}\geq\frac{\sigma}{2\sqrt{\beta}}=\frac{\sigma}{\sqrt{2\alpha}}.$ (119)
@@
Let $\beta=\alpha/2$. Now we construct hypotheses as following, which is more complex than those in additive contamination model.
@@
Construct $d$ hypotheses, and denote $p_{i}$ as the distribution of $\mathbf{X}$ under $i$-th hypothesis. If $Z=i$, then $\mathbf{X}$ follows $p_{i}$. $p_{i}$ is defined as follows. Given $Z=i$, $X_{i}=\sigma/(2\beta)$. Moreover, construct another auxiliary random variable $U$, such that $p_{i}(U=0)=1-\beta$, $p_{i}(U=1)=\beta$. Conditional on $U=0$, for all $j\neq i$, $p_{i}(X_{j}=0|U=0)=1-2\beta$, $p_{i}(X_{j}=\sigma/(2\sqrt{\beta})|U=0)=p_{i}(X_{j}=-\sigma/(2\sqrt{\beta})|U=0)=\beta$. Conditional on $U=1$, for all $j\neq i$, $p_{i}(X_{j}=\sigma/(2\beta)|U=1)=p_{i}(X_{j}=-\sigma/(2\beta)|U=1)=\beta$. Conditional on $Z=i$ and $U$, $X_{j}$ are independent for $j\in[d]\setminus\{i\}$.
Now the mean of $p_{i}$ is
$\mu_{i}=\frac{\sigma}{2\beta}\mathbf{e}_{i}.$ (122)
Under $Z=i$, for $j\neq i$, the variance in $j$-th dimension is
$\operatorname{Var}_{i}[X_{j}]=2\beta^{2}\left(\frac{\sigma}{2\beta}\right)^{2}+2\beta(1-\beta)\left(\frac{\sigma}{2\sqrt{\beta}}\right)^{2}\leq\sigma^{2},$ (123)
hence Assumption 2 is satisfied.
Construct $p_{Y}$, the distribution of $Y$ as following: $p_{Y}(Y_{j}=0)=1-2\beta$, $p_{Y}(Y_{j}=\sigma/(2\beta))=p_{Y}(Y_{j}=-\sigma/(2\beta))=\beta$. Then the total variation distance between $p_{i}$ and $p_{Y}$ is $\mathbb{TV}(p_{i},p_{Y})$. Hence, for any $i\in[d]$, $\mathbf{X}$ follows $p_{i}$, and the attacker can ensure that $\mathbf{Y}$ follows $p_{Y}$ and $\mathbf{Y}=\mathbf{X}$ with probability $\beta$. Note that now for $i,j\in[d]$, $i\neq j$, $\|\mu_{i}-\mu_{j}\|_{2}\geq\sigma/(\sqrt{2}\beta)$. Let $\beta=\alpha/2$, and follow similar steps in additive contamination model, with probability at least $1/2-\exp[-(\ln 2-1/2)N\alpha]$,
$\|\hat{\mu}-\mu^{*}\|_{2}\geq\frac{\sigma}{\sqrt{2}\alpha}.$ (124)
The proof is complete.
Appendix D Proof of Theorem 4
@@
From the condition (29), $t\leq\sigma$ holds. Then
@@
$=$ $\|\mathbf{U}_{p}^{T}(\mu(A)-\nabla F(\mathbf{w}))\|_{2}^{2}.$
@@
$\text{P}\left(\|\mathbf{P}(\mu_{w}(A)-\nabla F(\mathbf{w}))\|_{2}^{2}>\frac{2p}{N_{A}}\ln\frac{p}{\delta_{0}}\sigma^{2}\right)$ $=$ $\text{P}\left(\|\mathbf{U}_{p}^{T}(\mu(A)-\nabla F(\mathbf{w}))\|_{2}^{2}>\frac{2p}{N_{A}}\ln\frac{p}{\delta_{0}}\sigma^{2}\right)$ (133)
@@
Recall that there are $T$ iterations in total. Therefore, we get the following union bound. With probability at least $1-T\delta_{0}-Te^{-\frac{1}{64}\alpha m}-4Tme^{-\frac{1}{16}\lambda_{c}p\alpha^{2}m^{\frac{1}{3}}\epsilon^{2}}$, for $t=1,\ldots,T$, (140) holds for $\mathbf{w}_{1},\ldots,\mathbf{w}_{T}$. Let $\delta_{0}=\delta/T$. Then with probability at least $1-\delta-Te^{-\frac{1}{64}\alpha m}-4mTe^{-\frac{1}{16}\lambda_{c}p\alpha^{2}N^{\frac{1}{3}}\epsilon^{2}}$, for $t=1,\ldots,T$,
@@
in which $\rho$ and $\Delta$ are defined in (31) and (32). Then it is straightforward to show by induction that for $t=1,\ldots,T$,

### v3

@@ Appendix A Appendix
$\text{P}(\hat{Z}\neq Z)$ $\overset{(a)}{\geq}$ $\frac{H(Z|\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m},\mathbf{X}_{0})-\ln 2}{\ln(d-1)}$ (122)
$=$ $\frac{H(Z)-I(Z;\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m},\mathbf{X}_{0})-\ln 2}{\ln(d-1)}$
$\geq$ $1-\frac{I(Z;\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m},\mathbf{X}_{0})+\ln 2}{\ln d}$
$\overset{(b)}{=}$ $1-\frac{I(Z;\mathbf{X}_{0})+\ln 2}{\ln d}.$
(a) holds since $Z\rightarrow\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m},\mathbf{X}_{0}\rightarrow\hat{Z}$ is a Markov chain. (b) holds the distribution of $\mathbf{Y}$ does not depend on $Z$. The mutual information can be bounded by
$I(Z;\mathbf{X}_{0})=I\left(Z;\frac{1}{k}\sum_{i=1}^{k}\mathbf{U}_{i}\right)\leq I(Z;\mathbf{U}_{1},\ldots,\mathbf{U}_{k})=k(H(\mathbf{U}_{i})-H(\mathbf{U}_{i}|Z)),$ (123)
For any $j$, given $Z=j$, $U_{ij}$ is fixed. Therefore
$H(\mathbf{U}_{i})-H(\mathbf{U}_{i}|Z=j)$ $\leq$ $-2\beta\ln\beta-(1-2\beta)\ln(1-2\beta)$ (124)
@@
$\text{P}(\hat{Z}\neq Z)\geq 1-\frac{2k\beta\left(\ln\frac{1}{\beta}+1\right)+\ln 2}{\ln d}.$ (125)
Recall (111) yields
$k\leq\frac{\ln\frac{d}{4}}{4\beta\left(\ln\frac{1}{\beta}+1\right)}.$ (126)
Hence $\text{P}(\hat{Z}\neq Z)\geq 1/2$.
Recall (120). For $\in[d], iji\neq j$, $\|\mu_{i}-\mu_{j}\|_{2}\geq\sigma/\sqrt{\beta}$. Therefore, if $\hat{Z}\neq Z$, then
$\|\hat{\mu}-\mu^{*}\|_{2}$ $\overset{(a)}{=}$ $\|\hat{\mu}-\mu_{Z}\|_{2}$ (127)
@@
$\overset{(c)}{\geq}$ $\frac{\sigma}{2\sqrt{n\beta}}.$
For (a), note that given $Z=j$, the real mean of $\mathbf{X}$ is $\mu^{*}=mu_{i}$. (b) uses (121), which yields $\|\hat{\mu}-\mu_{\hat{Z}}\|_{2}\leq\|\hat{\mu}-\mu_{Z}\|_{2}$. (c) uses (120), which yields $\|\mu_{j}-\mu_{l}\|_{2}=\sigma/\sqrt{n\beta}$ for $j\neq l$.
From the above analysis, we conclude that if (126) holds, then with probability at least $1/2$, $\|\hat{\mu}-\mu^{*}\|_{2}\geq\sigma/(2\sqrt{\beta})$. Taking $\beta=\alpha/2$. From Lemma 11, with probability at least $1/2-\exp[-(\ln 2-1/2)m\alpha]$,
$\|\hat{\mu}-\mu^{*}\|_{2}\geq\frac{\sigma}{2\sqrt{n\beta}}=\frac{\sigma}{\sqrt{2n\alpha}}.$ (128)
The above derivation holds for $N_{A}\geq n$. Similar bound can be derived for $N_{A}<n$:
$\|\hat{\mu}-\mu^{*}\|_{2}\geq=\frac{\sigma}{\sqrt{2N_{A}\alpha}}.$ (129)
Therefore (112) holds.
@@
Let $\beta=\alpha/2$. Now we construct hypotheses as following, which is more complex than those in additive contamination model. For simplicity, now we only show the case with $N_{A}\geq n$.
@@
Construct $d$ hypotheses, and denote $p_{j}$ as the distribution of $\mathbf{X}$ under $i$-th hypothesis. If $Z=j$, then $\mathbf{X}$ follows $p_{j}$. $p_{j}$ is defined as follows. Given $Z=j$, for $i\in[m]$, $X_{ij}=\sigma/(2\sqrt{n}\beta)$. Moreover, construct another auxiliary random variable $R$, such that $p_{j}(R=0)=1-\beta$, $p_{j}(R=1)=\beta$. Conditional on $R=0$, for all $j\neq i$, $p_{j}(X_{j}=0|R=0)=1-2\beta$, $p_{j}(X_{j}=\sigma/(2\sqrt{n\beta})|R=0)=p_{j}(X_{j}=-\sigma/(2\sqrt{n\beta})|R=0)=\beta$. Conditional on $R=1$, for all $j\neq i$, $p_{j}(X_{j}=\sigma/(2\sqrt{n}\beta)|R=1)=p_{j}(X_{j}=-\sigma/(2\sqrt{n}\beta)|R=1)=\beta$. Conditional on $Z=j$ and $U$, $X_{j}$ are independent for $j\in[d]\setminus\{i\}$.
Now the mean of $p_{j}$ is
$\mu_{i}=\frac{\sigma}{2\sqrt{n}\beta}\mathbf{e}_{i}.$ (132)
Under $Z=j$, for $j\neq i$, the variance in $j$-th dimension is
$\operatorname{Var}_{i}[X_{j}]=2\beta^{2}\left(\frac{\sigma}{2\sqrt{n}\beta}\right)^{2}+2\beta(1-\beta)\left(\frac{\sigma}{2\sqrt{n\beta}}\right)^{2}\leq\frac{\sigma^{2}}{n},$ (133)
hence (108) is satisfied.
Construct $p_{Y}$, the distribution of $Y$ as following: $p_{Y}(Y_{j}=0)=1-2\sqrt{n}\beta$, $p_{Y}(Y_{j}=\sigma/(2\sqrt{n}\beta))=p_{Y}(Y_{j}=-\sigma/(2\sqrt{n}\beta))=\beta$. Then the total variation distance between $p_{j}$ and $p_{Y}$ is $\mathbb{TV}(p_{j},p_{Y})$. Hence, for any $i\in[d]$, $\mathbf{X}$ follows $p_{j}$, and the attacker can ensure that $\mathbf{Y}$ follows $p_{Y}$ and $\mathbf{Y}=\mathbf{X}$ with probability $\beta$. Note that now for $\in[d], iji\neq j$, $\|\mu_{i}-\mu_{j}\|_{2}\geq\sigma/(\sqrt{2n}\beta)$. Let $\beta=\alpha/2$, and follow similar steps in additive contamination model, with probability at least $1/2-\exp[-(\ln 2-1/2)N\alpha]$,
$\|\hat{\mu}-\mu^{*}\|_{2}\geq\frac{\sigma}{\sqrt{2n}\alpha}.$ (134)
The case with $N_{A}<n$ can be proved similarly. The proof of Theorem 3 is complete.
Proof of Theorem 4
@@
From the condition (condition (5) in Theorem 4)
$N_{A}\geq 2\ln\frac{mT}{\delta},$ (141)
$t\leq\sigma$ holds. Then
@@
$=$ $\|\mathbf{U}_{p}^{T}(\mathbf{X}_{0}-\nabla F(\mathbf{w}))\|_{2}^{2}.$
@@
$\text{P}\left(\|\mathbf{P}(\mu_{w}(A)-\nabla F(\mathbf{w}))\|_{2}^{2}>\frac{2p}{N_{A}}\ln\frac{p}{\delta_{0}}\sigma^{2}\right)$ $=$ $\text{P}\left(\|\mathbf{U}_{p}^{T}(\mathbf{X}_{0}-\nabla F(\mathbf{w}))\|_{2}^{2}>\frac{2p}{N_{A}}\ln\frac{p}{\delta_{0}}\sigma^{2}\right)$ (144)
@@
Recall that there are $T$ iterations in total. Therefore, we get the following union bound. With probability at least $1-T\delta_{0}-Te^{-\frac{1}{64}\alpha m}-4Tme^{-\frac{1}{16}\lambda_{c}p\alpha^{2}m^{\frac{1}{3}}\epsilon^{2}}$, for $t=1,\ldots,T$, (151) holds for $\mathbf{w}_{1},\ldots,\mathbf{w}_{T}$. Let $\delta_{0}=\delta/T$. Then with probability at least $1-\delta-Te^{-\frac{1}{64}\alpha m}-4mTe^{-\frac{1}{16}\lambda_{c}p\alpha^{2}m^{\frac{1}{3}}\epsilon^{2}}$, for $t=1,\ldots,T$,
@@
in which $\rho$ and $\Delta$ are defined as $\rho=\eta\mu/2$ and
Then it is straightforward to show by induction that for $t=1,\ldots,T$,